40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 48 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  243   Wed Jul 28 22:17:15 2010 FrankNotesDAQsome connections to P4-3113A (J2 and J3)

- personal notes -

 

J2  63/64  CH31  PSL-ISS_AOMRF

J2  45/46  CH22  PSL-FSS_LODET

J2  49/50  CH24  PSL-FSS_FAST

J2  51/52  CH25  PSL-FSS_PCDRIVE

J2  43/44  CH21  PSL-FSS_RFPDDC

J2  53/54  CH26  PSL-FSS_RCTRANSPD

J2  47/48  CH23  PSL-FSS_PCDET

 

J3  ?   CH34  PSL-PMC_TRANSPD

  242   Wed Jul 28 22:12:41 2010 FrankNotesDAQold connections - J4 / J5

- personal notes -

VMIC3123 (16bit)

J-4:

12/25 : FSS_RMTEMP

6/19 : FSS_MINCOMEAS

 

J-5:

2/14 : FSS_RCTEMP

  241   Wed Jul 28 22:11:11 2010 FrankNotesDAQold channel connections - blocks TB2 and TB4

- personal notes -

current cross-connect connections

PMC_RFPDDC:
block TB4
1 - LO
2 - HI
connected to ?

------------------------------------------------------------------------------

FSS_RFPD_DC:
block TB2
1 - LO
2 - HI
connected to J2-3113A  43/44 (CH21)

------------------------------------------------------------------------------

FSS_RCTRANSPD:
block TB2
5 - LO
6 - HI
connected to J2-3113A  53/54 (CH26)

------------------------------------------------------------------------------

PMC_TRANSPD:
block TB4
3 - LO
4 - HI
connected to J3-3113A   (CH34) (?)

 

 

  240   Wed Jul 28 21:53:10 2010 FrankNotesDAQconfusing cross-connect panel connections

the documents we have in the PSL lab about the connection made in the PSL rack cross-connect panel are not consistent with the current wiring.
A couple of channels are labeled to be connected to the VMIC3123 card (16bit) , but are connected to the VMIC3113A card (12bit).
One of these channels is  PSL-FSS_RCTRANSPD. I will reconfigure some of those channels.
Documentation will be provided in the elog and changes marked in the existing, printed version of cross-connect panel document.

  239   Wed Jul 28 19:47:58 2010 MeganNotes Slow actuator calibration

For future people searching in the elog, calibration for the slow actuator is 1182 MHz/V. Or for the temperature actuator. Whichever one decides to search in that moment.

  238   Wed Jul 28 19:46:46 2010 MeganDailyProgress PSL Progress!

The PMC is still locked (except when we do something to unlock it) and both the analyzer cavities and reference cavities have the ability to be locked. We have not locked them simultaneously yet but will be working on that so we can get a beat. We connected the VCO driver to the AOM for the analyzer cavity and got it locked.

A preliminary calibration for the photodetectors in the path of the transmitted light from the reference cavity has been calculated, although changes in the next day or two might change it, so it has not been put into the computer.

The table by the cavities is now very pretty . Until we fill all the space we just cleared with everything we need to beat the two beams together!

Preparations for that will be started tomorrow by putting the correct optics in the path of the transmitted light of the analyzer cavity and better aligning those in the path of the reference cavity's transmitted light.

Also, as an extra note, last week we calibrated some inputs, so we no longer have anti-energy reflected from the PMC! In fact, at this moment, we have 1.4mW. Success is still 100 and strength is down to 44! Guess it's time for dinner.

  237   Mon Jul 26 22:14:39 2010 MeganDailyProgress Daily Update

Today I aligned the AOM so that I could align the analyzer cavity to lock the cavity. We already aligned and locked the PMC last week so we had a beam to work with. The AOM is now aligned and the beam coming from it is aligned into the analyzer cavity. We scanned the cavity to find the mode we want and currently the output of the photocell is being mixed with the local oscillator to find the error signal. It's not quite how it should be right now, so we will look at it tomorrow to figure out what is wrong. Then we can connect it to the PDH servo and lock the cavity!

I also did some cable-management. They are now neatly organized and held in place by a large number of zip ties so they don't get in the way of everything else! The cameras are now connected so that we can see the beam from the PMC, analyzer cavity, and reference cavity.

  236   Mon Jul 26 20:43:30 2010 FrankSummaryComputersnew workstation

i got an old workstation and have converted it into a new linux workstation for the PSL lab.
It's like one of the usual workstation but with only a single screen due to space limitations.

Name: WS5
IP-address: 10.0.0.25

users are root and controls. passwords the usual ones...

  235   Wed Jul 21 10:39:10 2010 FrankDailyProgressComputersMaster MEDM screen for RefCav experiment

i don't know the units for the strength, but it's showing how good the reception of the signal is.I can't be dBm or something like that, so i don't know what units they use for that kind of quality signal.

The success rate is basically percent. Same situation here, there is no explanation for this signal, but it goes from 0 to 100 so it's percent.
If the value drops below 100 that basically means that the sensor is close to the range limit or the batteries are empty and so the communication is bad and we are loosing some data points...

the reason for some funny numbers like getting energy out of the PMC is that most of those channels are not connected or not calibrated yet. The PMC refl power is an open input

Quote:

While I guess it's a nice idea to put a Success = 100 meter on all of our MEDM screens for motivational purposes, what are the units on this one / the meaning of it?  Everything else makes sense (except maybe for the strength....are strength and success related?), but the success is confusing.

Quote:

i prepared a master screen for the refcav experiment.
This main screen contains all important numbers to figure out what the status of the experiment is.
It also contains several stripcharts which monitor important values like temperatures of the cavities, transmitted power levels etc over time.
It also has several buttons which link to the individual sub-screens which already exist for our subsystems.

 

 

  234   Mon Jul 19 18:21:50 2010 taraSummaryNoiseBudgetLaser intensity noise & pointing

I checked the noise budget code psl_refcav_sio2_300K.m on Laser intensity noise.

The value for df/dp, which is intensity dependence of the cavity resonance frequency, from Alnis et al. is 70 Hz/uW (page 6, just before section 3)

for the light coupled into the cavity. I'm not sure where the original value, 25 Hz/uW, in the code came from. I'll ask Rana about this.

 df/dp already incorporates both effects from photo elastic and photo refractive noise in the coatings + substrate,

so we don't have to worry about dn/dT, or

thermal expansion coeff of the material.

Their coating materials are SiO2, Ta2O5, 38 layers. Our coating materials are similar, but only 17 layers thick.

So we should be able to rescale their df/dp for our cavity.

                                        | Alnis etal|    PSL   |

cavity length [mm]        |  77.5      |  203.5  |

Finesse                          | 4e5         |  1e4     |

wavelength [nm]           |  972        |   1064  |    

If we have dp (the power fluctuation of the coupled power, coupling efficiency * input power * RIN,),df can be calculated.

However, df should depend on the cavity length, and the nominal frequency of the beam. 

df ~  f dL / L, where f = nominal frequency, L = cavity length.

And dP should depend on the circulating power inside the cavity which truly tells the value of absorbed power on the coatings.

dp ~ Pcirculating/ Finesse.

Thus, our df/dp should be rescaled to

70e6 * (203.5  /  77.5 ) * (1e4 / 4e5) = 4.6e6 Hz/W.

=========================================================

About Pointing noise. The instability of laser pointing will cause power fluctuation coupled into the cavity.

  This power fluctuation will induce frequency noise. If we know

how the beam moves, power coupled into the cavity and fluctuation can be calculated.

I'm thinking about using 2 QPDs to measure the beam position. A beam splitter will split the beam between 2 QPDs, the difference between

two readouts should tell us about angular and tranlational motion, and the power will be

P = Po Exp [ - (dx/w0)2 - (dtheta/ div angle)2]

where dx = translational motion

w0 = beam's waist

dtheta = angular motion

div angle = beam divergence angle.

  233   Mon Jul 19 03:39:51 2010 RanaDailyProgressComputersMaster MEDM screen for RefCav experiment

Leave Frank alone!  His strength is already down to 47 and his PMC has -2mW of anti-energy coming out of it.

  232   Sun Jul 18 23:59:52 2010 JenneDailyProgressComputersMaster MEDM screen for RefCav experiment

While I guess it's a nice idea to put a Success = 100 meter on all of our MEDM screens for motivational purposes, what are the units on this one / the meaning of it?  Everything else makes sense (except maybe for the strength....are strength and success related?), but the success is confusing.

Quote:

i prepared a master screen for the refcav experiment.
This main screen contains all important numbers to figure out what the status of the experiment is.
It also contains several stripcharts which monitor important values like temperatures of the cavities, transmitted power levels etc over time.
It also has several buttons which link to the individual sub-screens which already exist for our subsystems.

 

  231   Sun Jul 18 18:32:08 2010 FrankDailyProgressComputersMaster MEDM screen for RefCav experiment

i prepared a master screen for the refcav experiment.
This main screen contains all important numbers to figure out what the status of the experiment is.
It also contains several stripcharts which monitor important values like temperatures of the cavities, transmitted power levels etc over time.
It also has several buttons which link to the individual sub-screens which already exist for our subsystems.

medmScreen.png

  230   Fri Jul 16 16:58:02 2010 FrankNotesRefCavlatest values for locking refcav

 - personal notes -

PMC voltage : 212V
PMC transPD : 0.291V

slow actuator : 0.1500V
RCTransPD : 0.796V

  229   Fri Jul 16 15:34:51 2010 MeganDailyProgressVCOBusby box

Quote:

there are 0 Ohm resistors, but you can also make a simple short between those two pads

i don't understand the third graph.  Looks like the blue data is the frequency noise calculated from the electronic noise with your 680KHz/V coefficient. The green one looks like phase noise, so i think you have to convert one of those into other units

Quote:

I replaced the batteries of the Busby box today to see if that reduced noise. It did not. Measuring the instrument noise and TP7 noise gave exactly what I got yesterday with the old batteries. I subtracted the instrument noise to plot against the model again and got basically the same graph as yesterday.

I also tried comparing the phase noise coming from the electronic noise against the total noise of the VCO. The total noise of the VCO was apparently below that of the electronic noise. I'm wondering if it could have been the fault of all the ground loops, since removing almost every cable changed the graph of the electronic noise so much. So at some point soon I would like to re-connect the entire VCO and measure the phase noise again, making sure I remove every unnecessary cable. This part confuses me though - all the schematics label the connecting resistor as 0, so I'm not sure if there is a 0 resistor or if it's a value that I don't know?

 

 Looks like I made the third graph without really thinking about what I was doing. The y-axis is labeled completely incorrectly, the blue data is frequency noise calculated from the electronic noise, and the green data is the phase noise from the VCO. So, after thinking about what I was doing, I converted the electronic noise into phase noise and graphed them together (with correct axis labels). They look more similar now, but I would still like to measure the VCO without the extra cables. I'll try connecting the VCO and measure again.

Attachment 1: PhaseNoiseCompared.png
PhaseNoiseCompared.png
  228   Fri Jul 16 15:18:55 2010 FrankDailyProgressVCOBusby box

there are 0 Ohm resistors, but you can also make a simple short between those two pads

i don't understand the third graph.  Looks like the blue data is the frequency noise calculated from the electronic noise with your 680KHz/V coefficient. The green one looks like phase noise, so i think you have to convert one of those into other units

Quote:

I replaced the batteries of the Busby box today to see if that reduced noise. It did not. Measuring the instrument noise and TP7 noise gave exactly what I got yesterday with the old batteries. I subtracted the instrument noise to plot against the model again and got basically the same graph as yesterday.

I also tried comparing the phase noise coming from the electronic noise against the total noise of the VCO. The total noise of the VCO was apparently below that of the electronic noise. I'm wondering if it could have been the fault of all the ground loops, since removing almost every cable changed the graph of the electronic noise so much. So at some point soon I would like to re-connect the entire VCO and measure the phase noise again, making sure I remove every unnecessary cable. This part confuses me though - all the schematics label the connecting resistor as 0, so I'm not sure if there is a 0 resistor or if it's a value that I don't know?

 

  227   Fri Jul 16 14:19:02 2010 MeganDailyProgressVCOBusby box

I replaced the batteries of the Busby box today to see if that reduced noise. It did not. Measuring the instrument noise and TP7 noise gave exactly what I got yesterday with the old batteries. I subtracted the instrument noise to plot against the model again and got basically the same graph as yesterday.

I also tried comparing the phase noise coming from the electronic noise against the total noise of the VCO. The total noise of the VCO was apparently below that of the electronic noise. I'm wondering if it could have been the fault of all the ground loops, since removing almost every cable changed the graph of the electronic noise so much. So at some point soon I would like to re-connect the entire VCO and measure the phase noise again, making sure I remove every unnecessary cable. This part confuses me though - all the schematics label the connecting resistor as 0, so I'm not sure if there is a 0 resistor or if it's a value that I don't know?

Attachment 1: NewBusbyNoise.png
NewBusbyNoise.png
Attachment 2: BusbyModelNoise.png
BusbyModelNoise.png
Attachment 3: PhaseNoiseCompared.png
PhaseNoiseCompared.png
  226   Thu Jul 15 20:30:30 2010 FrankNotesPMCPMC scan

PMC scanned using slow actuator (NPRO temp)

pmc-scan.png

pmc-scan2.png

  225   Thu Jul 15 19:23:51 2010 FrankNotesPEMtemp sensor calibration using PT1000

platinum temp sensor (PT1000) used for calibration of AD590 on PSL table:

Callendar-Van Dusen Equation:
RT = R0(1+AT+BT2-100CT3+CT4)
RT = Resistance at temperature T (°C)
R0 = Resistance at 0°C
T = Temperature in °C

 

CALLENDAR-VAN DUSEN CONSTANTS
Alpha, a (°C-1) .003750 ± .00003 .003850 ± .0001
Delta, d (°C) 1.605 ± 0.009 1.4999 ± 0.007
Beta, b* (°C) 0.16 0.10863
A (°C-1) 3.81*10e-3 3.908*10e-3
B (°C-2) -6.02 *10e-7 -5.775*10e-7
C (°C-4)* -6.0*10e-12 -4.183*10e-12

*Both b =0 and C=0 for T>0°C

source:  http://content.honeywell.com/sensing/prodinfo/temperature/technical/c15_136.pdf
online calculator: http://www.isotech.co.uk/prtcalc-web.html

 

measured resistance: 1083.9Ohm = 21.55C

calibration offsets for chamber sensors:
S1: +0.75K
S2: 0
S3: -0.11
S4: +0.88

room temp sensors:

ACAV_AMB1 : +0.2
ACAV_AMB2:  -0.3

  224   Thu Jul 15 19:04:45 2010 FrankDailyProgressVCOVCO Noise Update

looks quite good.

You can take the two new batteries out of the FET preamp box if you can't find more (the ones we replaced last week).

i had a quick look onto your new data. If you subtract the busby box noise floor from the new data you are almost there...

Quote:

I re-measured the noise at TP7 using the Busby Low Noise Box and measured the instrument noise of the box - the noise is lower at high frequencies, but much higher at low frequencies. There is a possibility some of this came from old batteries - they measured 7.4 V, but we were 2 short in 40m so I couldn't replace them. I will try to get new batteries tomorrow and see if that affects it at all.

I also changed the couple of things that were wrong with the model and have the revised model as well as its resultant graph attached.

I subtracted the noise of the other amplifier from the previous measurement of the noise at TP7 and compared it with the new model and they line up almost perfectly except around 10 Hz, where they're a little off.

 

  223   Thu Jul 15 18:18:49 2010 MeganDailyProgressVCOVCO Noise Update

I re-measured the noise at TP7 using the Busby Low Noise Box and measured the instrument noise of the box - the noise is lower at high frequencies, but much higher at low frequencies. There is a possibility some of this came from old batteries - they measured 7.4 V, but we were 2 short in 40m so I couldn't replace them. I will try to get new batteries tomorrow and see if that affects it at all.

I also changed the couple of things that were wrong with the model and have the revised model as well as its resultant graph attached.

I subtracted the noise of the other amplifier from the previous measurement of the noise at TP7 and compared it with the new model and they line up almost perfectly except around 10 Hz, where they're a little off.

Attachment 1: BusbyNoise.png
BusbyNoise.png
Attachment 2: CorrectModelMeasured.png
CorrectModelMeasured.png
Attachment 3: NoiseAll.fil
# All noise
  # Naming convention due to repeat pin numbers
    # All points not at a test point labelled with n number of the next test pin _ pin number ie n6_3
    # All test points labelled by TP(number) ie TP4
    # All points with no pin number labelled with preceding pin number, following pin number ie n7_26 (between 2 and 6)

# Wide SW to amplifier

  # Up to TP6

... 82 more lines ...
Attachment 4: CompleteModel.png
CompleteModel.png
  222   Thu Jul 15 16:34:11 2010 MeganNotesVCOBusby Low Noise Box

I brought the Busby low noise box back to 40m to take more measurements.

  221   Thu Jul 15 15:46:20 2010 FrankNotesVCOMatching Model and Measurement

we replaced U9 by an OP27 instead of AD797. The replacement AD797 and AD829 were oscillating at 12MHz and larger.
So if you replace U9 by an OP27in your model the noise will be slightly higher at high frequencies.

Did you subtract the instrumentation noise from your measurement @TP7?
If not then do it.
As far as i remember the noise level of the DAQ/preamp is/was about 4nV/sqrt(Hz). I guess the simulation will then match your measurement

Quote:

Thanks, I modeled the LM336 and get almost the same measured curve at low frequencies. I'm on my way over to 40m in a matter of minutes to remeasure the noise with a lower noise amplifier to see if the entire curve can fit. I also have the model attached - the strange graphing frequencies were to match the ones I got with the measured data. I am also going to double-check the opamps used in the circuit currently. I already verified the resistor values - only R18 was incorrect (1k instead of 390).

 

  220   Thu Jul 15 15:08:46 2010 FrankNotesPMCcurrent values to lock system

personal notes to lock system remotely:

for PMC:
PMC voltage 203.5V --> DC offset -3.258
slow DC value : 0.15
PMC transmission value: 0.291V

  219   Thu Jul 15 14:13:01 2010 MeganNotesVCOMatching Model and Measurement

Thanks, I modeled the LM336 and get almost the same measured curve at low frequencies. I'm on my way over to 40m in a matter of minutes to remeasure the noise with a lower noise amplifier to see if the entire curve can fit. I also have the model attached - the strange graphing frequencies were to match the ones I got with the measured data. I am also going to double-check the opamps used in the circuit currently. I already verified the resistor values - only R18 was incorrect (1k instead of 390).

Attachment 1: ModelMeasuredNoise.png
ModelMeasuredNoise.png
Attachment 2: NoiseAll.fil
# All noise
  # Naming convention due to repeat pin numbers
    # All points not at a test point labelled with n number of the next test pin _ pin number ie n6_3
    # All test points labelled by TP(number) ie TP4
    # All points with no pin number labelled with preceding pin number, following pin number ie n7_26 (between 2 and 6)

# Wide SW to amplifier

  # Up to TP6

... 82 more lines ...
  218   Thu Jul 15 13:47:04 2010 FrankNotesVCOLM 336 Noise

you can model voltage references like this:

use an ideal opamp and add the parameters for the voltage noise manually: un=flat noise level, uc=1/f corner frequency

op LM336 op0 nii no no un=205e-9 uc=80   # simulate noise of LM336-5 reference

can you post the whole vco model plz. I would like to check the whole model. you might use the wrong opamps as your schematic is not the one for the vco you are debugging...

Quote:

I had forgotten to include the noise of the LM336 in the noise model because it cannot be easily modeled by LISO. So I thought I would look up its noise to see if I could make an approximation just to see if it will affect the noise levels coming from the model. I found this on a data sheet and will try to find a way to include it in the model.

 

  217   Thu Jul 15 13:46:48 2010 MeganNotesVCOTest Points Modeled and Compared

I modeled the noise at each test point (ignoring the LM366) and compared with the current measured data. As we already know, some of the measurements are limited by instrumental noise, so I will update these graphs as soon as I have data not limited by amplifier noise, but the current graphs support the idea that the noise of the LM366 might account for some of the difference since the measured data at TP6 (the only one independent of the LM366) coincides more closely with the model. We will see if that continues or even improves after I take new, not noise limited, data.

Attachment 1: TPsCompared.png
TPsCompared.png
  216   Thu Jul 15 11:43:36 2010 MeganSummaryVCOMeasured and Modeled Noise with Instrumentation Noise

I created a very rough model of the LM336 (emailed to Frank and Jan to try to improve) to see if the noise of the LM336 might explain the noise at lower frequencies. I think from this rough sketch that might be the case, but again, it was a very rough model. This shows the new model with the instrumentation noise and measured noise. The instrumentation noise probably explains the higher frequency discrepancy - I will test that later today - and there's a chance that the LM336 explains the lower frequency discrepancy.

Attachment 1: MeasModelInstNoise.png
MeasModelInstNoise.png
  215   Thu Jul 15 10:38:03 2010 MeganNotesVCOLM 336 Noise

I had forgotten to include the noise of the LM336 in the noise model because it cannot be easily modeled by LISO. So I thought I would look up its noise to see if I could make an approximation just to see if it will affect the noise levels coming from the model. I found this on a data sheet and will try to find a way to include it in the model.

Attachment 1: LM336Noise.PNG
LM336Noise.PNG
  214   Wed Jul 14 19:07:17 2010 MeganDailyProgressVCOVCO Update

I replace R3 and R7 with thin film resistors (according to LISO they have the greatest contribution of the resistors at lower frequencies) and re-measured the electronic noise of the driver. It was almost identical to the noise I previously measured, and so for now am not going to replace the other resistors because it does not seem to have a large effect.

I have been playing around with the LISO model more and will hopefully have an easy-to-use version to upload soon. I'm going to work with more of the data that I have to see what it tells us and will post various graphs tomorrow.

  213   Tue Jul 13 18:40:56 2010 MeganSummaryVCOComparison of Modeled and Measured VCO Noise

I modeled the VCO using LISO to find a model of the electronic noise entering the actual VCO. I will double-check all the resistor and capacitor values with the actual VCO to make sure I'm modeling what I actually measured and post the file I used. I also converted the electronic noise into frequency noise of the VCO assuming the conversion factor of 680 kHz/V (found by measuring the frequency response around 80MHz - previous post).

I will work on replacing resistors tomorrow, after looking at my model to see which have the greatest impact on noise, and see if that decreases the measured electronic noise.

Attachment 1: Electronic_and_Frequency_Noise.png
Electronic_and_Frequency_Noise.png
  212   Tue Jul 13 14:46:23 2010 taraNotesFSSschematic for FSS servo

The current FSS servo we are using is labeled D980536 Rev. C on the PCB, but the schematic that matches the PCB more is D980536-D (there are minor corrections).

The schematic is posted below.

I'm checking If the board will match the schematic or not. Then its transfer function (TF) will be calculated by simulink, and compared with the result from measurement. 

The correction will be added soon.

 

Attachment 1: D980536-D.pdf
D980536-D.pdf
  211   Tue Jul 13 11:35:12 2010 MeganSummaryVCOVCO Frequency Response

We measured the frequency response of the VCO and fit it with a linear curve to give an idea of the frequency response around 80MHz. This shows that the general trend is not linear, but can be approximated within 1V applied to the VCO.

Attachment 1: VCOFreqResponse.png
VCOFreqResponse.png
  210   Mon Jul 12 14:32:49 2010 taracSummaryRefCavRCTRANSPD RIN

Before, when we tried to lock the cavity, it seemed that the gain was too high, and the power output

from RefCav(C3:PSL-FSS_RCTRANSPD) fluctuated when both PC and Fast feedback were connected, and the data was measured when only Fast feedback was on.

So I try to change RF power to reduce the modulation index (hence, the gain.)

The value before was 6.5V, This time I tried 5.5 and 5.8 V.

At each RF power setting, I lock the cavity with

1) Fast feedback alone, and

2) both Fast and PC feedback

and measure the transmitted beam's RIN. The gain for each setup was readjust (see the values for each setup below)

so that the transmitted beam power was most stable as seen on an oscilloscope, then the data was taken.

The results are plotted below. It turns out that

when the PC feedback is connect, RIN becomes noisier. I don't no where 

it comes from. It might be that the setting still has too high gain, or RFAM from broadband EOM due to misalignment,

or broken opamp in FSS PC path.

 


 I'm still not sure why there are jumps at 100 Hz and 1k Hz. The sr785 is set at 4 different spans

which are, 0-100Hz, 0-1kHz,0-10kHz, and 0- 100kHz. Each span contains 800 FFT lines. The auto range is on.

 

The setting are [this will be updated soon]

1)RF5.5

 Fast feedback: on

 PC feedback: off

 CG

 FG

 

2) RF5.5

  Fast: on

  PC: on

 CG:

 FG:

 

3 RF 5.8

Fast: on

  PC: off

 CG:

 FG:

 

 

4 RF 5.8

Fast: on

  PC: on

 CG:

 FG:

 

 

 

 

 

 

Attachment 1: RIN.png
RIN.png
  209   Mon Jul 12 11:49:17 2010 MeganSummaryVCONoise of 2023B and Electronic Noise of VCO (correct graphs)

The previous graphs are wrong. Here are the correct ones and the electronic noise is plotted compared to the total noise of the VCO.

Attachment 1: NoiseBothTypes.png
NoiseBothTypes.png
Attachment 2: IndividualNoiseBothTypes.png
IndividualNoiseBothTypes.png
Attachment 3: VCOElectronicNoise.png
VCOElectronicNoise.png
  208   Mon Jul 12 11:04:16 2010 MeganSummaryVCONoise of 2023B and Electronic Noise of VCO

I got the measurements locking the 2023B (from the PSL lab) to one of the 2023As in 40m. I have one graph showing the calibrated data compared to the calibrated data of one of the other 2023As locked to the one I used. Another has the individual phase noise of the 2023A and the 2023B (by removing the noise of the 2023A). The individual noise of the 2023B can be found by using the equation

S2023B = sqrt(S2023A2 - SCombined2)

The individual noise looks very close to that of the other Marconis (previous graphs show all 3 2023As are almost identical), so I'm not going to test the 2023B at all frequency combinations.

I also measured the electronic noise of the VCO, only accounting for the gain of 1000 in the feedback loop.

Attachment 1: NoiseBothTypes.png
NoiseBothTypes.png
Attachment 2: IndividualNoiseBothTypes.png
IndividualNoiseBothTypes.png
Attachment 3: ElectronicNoise.png
ElectronicNoise.png
  207   Sun Jul 11 00:52:14 2010 MeganSummaryVCOVCO Noise

I graphed the VCO phase noise with the previously measured data - the data I have seems to be a bit higher. I also converted the phase noise into frequency noise for both VCOs and the previous measurement. Just for reference, "VCO 1" is the one that we were measuring the noise of the different parts of the VCO and that has its lid off.  "VCO 2" is the one we brought over from the PSL lab that is currently still in one piece.

Attachment 1: PhaseNoisePreviousMeasure.png
PhaseNoisePreviousMeasure.png
Attachment 2: FrequencyNoisePreviousMeasure.png
FrequencyNoisePreviousMeasure.png
  206   Sat Jul 10 05:31:15 2010 ranaSummaryNoiseBudgetfrom the SVN today

I modified the photo-thermal part of the NB to use the Alnis et. al. number of 25 Hz / uW  (frequency shift per input power noise). I adjusted it for the ratio of the cavity lengths between theirs and ours, and also for the ratio of the Finesses (400000 : 10000).

Tara should now grab this code and replace my guess at the RC RIN with his real measurements and also add the previously measured RC/AC frequency noise to this plot.

The punchline is that the this thermo-refractive noise is significant and we must control the low frequency RIN.

Attachment 1: psl_refcav_sio2_300K.png
psl_refcav_sio2_300K.png
  205   Fri Jul 9 11:45:05 2010 MeganSummaryVCOPhase Noise of Marconis and VCO

I used the data for the phase noise of the marconis to calculate the individual phase noise of each marconi. The noise of the three are roughly the same. The noise increases with frequency and input range, with the maximum variation in our measurements about 1 order of magnitude. This shows the noise is fairly consistent, but does change with different frequencies used and different input ranges.

We also have some data for the phase noise of the VCO compared with Marconi 2 (with the feedback going to the Marconi). The VCO is currently much noisier than the Marconis, so hopefully we can reduce the noise by modifying the current VCO.

We used swept sine to find the gain of the feedback loop, with a UGF of roughly 2.8 KHz with a gain of 1. With a gain of 100, the UGF is roughly 280 kHz and a peak to peak voltage when unlocked of 438mV.

The gain (for calibration) with the poles and zeroes can be calculated as:

(4V/216)(pi/Vp-p)(1/1000)(0.03)(UGF)

where 4/216 is the resolution of the ADC, pi/Vp-p gives the number of radians per volt, 1000 is the amplification of the signal into the ADC, 0.03 is the first zero (to compensate for the amplifier, which diminishes the signal below 0.03 Hz), and the UGF is the other zero. Multiplication by the zeros ensures the signal above the zeros will not be affected by the transfer function.

As a note - when the signals are calibrated in the DTT, the calibrated data can be saved by exporting the trace, not the signal itself, by exporting win0_pad0_trace1 or whichever trace it is saved under. This removes a need to calibrate the original data outside the DTT.

Attachment 1: IndividualNoise_80MHz_160MHz_10Hz_100HzRanges.png
IndividualNoise_80MHz_160MHz_10Hz_100HzRanges.png
Attachment 2: Separate_Individual_Noise.png
Separate_Individual_Noise.png
Attachment 3: VCONoise_Gain1_100.png
VCONoise_Gain1_100.png
  204   Thu Jul 8 19:55:42 2010 taraNotesFSSErr signal pk-pk vs RF adj

 Peter helped us enable the slow loop for the laser driver. The current setup is

P: +0.0001

I:   0.0000  (this is too sensitive, when I increase the gain by the smallest step, 0.0001, the system is out of lock)

D: 0.0000   

 

Also, after align the beam to the RefCav again. the transmitted beam from RefCav is much more stable.

 Reflected beam's Voltage on the RFPD is less than 20 mV when locked, and more than 496 mV when out of lock.

This corresponds to ~ 96% Transmission.

I measured the power spectral density of RFTRANSPD,

this is posted and compared to result we got from 40m on fig2.

*****This is done with slow loop(temperature) and fast feedback (pzt)  enabled, but PC feedback disconnected, because the gain is too high already.

Common gain is -2.5 dB,

Fast Gain is 15 dB.

[updated]

The blue plot is the result from 40m, red and green are the results from PSL lab.

Green was measured while the HEPA filter above the table was on, Red was measured

when the HEPA filter was off. I measured 4 Frequency span for each plot, 100k, 10k ,1k and 100 Hz. 

The results from PSL (red and green plot) have jumps at 100 and 1 kHz

which I haven't figured out yet. The next plan is to reduce the gain,

so that PC feedback (faster than 100 Hz) can be enabled. 

Thus, the RIN above 100 Hz should be suppressed more.


To reduce the gain, the RF power (modulating index) must be decreased. I don't know the calibration yet,

so I measure the pk-pk value of the error signal vs the RF power adjustment and plot below (fig1.) This might be helpful

for determining the right value.

 

 

Attachment 1: FSSerr.png
FSSerr.png
Attachment 2: RIN_compare.png
RIN_compare.png
  203   Thu Jul 8 17:17:10 2010 MeganDailyProgressVCOUpdate on what we have!

Frank and I made connectors for the VCOs so we don't have to have so many cords connected inside the VCO box. We also measured one of the VCOs and established it is a lot noisier than our Marconis. We'll work on modifying it and see how that changes it. Meanwhile we'll test the other to see if it's at the same level or different.

Also, I've now learned how to correctly calibrate the phase noise so I will hopefully be able to go back and calibrate a lot of the data that I have so we have a better idea of what is actually going on.

  202   Thu Jul 8 15:20:24 2010 FrankNotesRefCavtemp sensor calibration

PT1000 value (chamber temp):

1083.9 Ohm

 

AD590 values:

RT:20.6
S1: 20.8
S2: 21.55
S3: 21.76
S4: 20.77

  201   Wed Jul 7 20:02:53 2010 taraDailyProgressComputersBeam output is more stable

 

1) I just found out that the pbs + 1/4 waveplate in front of RefCav is not well aligned, so the length,as seen by the beam, is not 1/4 wavelength.

    Hence, the polarization is not turned 90 degree after double passing, and

    The reflected beam can go back to the laser and causes the power fluctuation we saw before.

    When the beam is blocked anywhere before RefCav, the beam output from PMC is very stable.

    I adjusted the PBS's angle and reduced the reflected power. Now the input power to PMC can go up to 50 mW without any fluctuation. 

 

2) I re-aligned the beam into RefCav, with Frank's help on gain adjusting,

    the transmitted power seems to be more stable. RefCav transmitted power RIN is posted below. I'll post the comparison between result at 40m soon. From a quick glance, RIN from 40m is at least 2 order of magnitude below

our result.

 

3) The PID control for slow actuator is up. I was adjusting it, but the medm screen was frozen.

    I reset it, and set the PID control (only P-part). The current setting for Proportional control(C3:PSL-FSS_SLOWKP) is +0.41.

 

Attachment 1: PSL_RF_RIN.png
PSL_RF_RIN.png
  200   Wed Jul 7 16:08:40 2010 MeganSummaryVCOUpdate on VCO offset

I realized I was doing the calculations incorrectly - this should make more sense.

Attachment 1: VCO_Output_Input.png
VCO_Output_Input.png
  199   Wed Jul 7 15:15:59 2010 MeganDailyProgressVCOPhase Noise Measurements

I'm currently working on measuring the phase noise of the Marconis at 80 MHz - I will be moving on to 160 MHz soon. I'm also working on learning how to make the computer do what I want it to, but I should be done with the measurements and post graphs later today. Then depending how long the measurements take today, I'll start measuring our Marconi and then move on to the VCO tomorrow and should be able to modify the VCO by Friday. One of the main things that's slowing me down is getting comfortable with processing the data on the computer.

Also, I've been having problems getting the Marconis to lock at any feedback gain below 2000. I've been using that to stay consistent and get a good lock between the two, because with lower gain there was always a sneaky little sine wave making it through the feedback loop and into the locked signal. I've accounted for this in the calibrations I've been making, with a UGF of around 1000 Hz.

  198   Wed Jul 7 10:44:39 2010 MeganSummaryVCOVCO Output vs Input

I redid one of the graphs for the output of the VCO vs the input voltage to make sure I didn't miscalculate. I got the same graph when accounting for the 19dBm attenuation at the output of the VCO. I also made a graph of the output RMS voltage ignoring the attenuation - the total output ranged from 13mV to 918mV, while the voltage added by taking the attenuation into account is 1.9929V.

I used the equation N dBm = 10*log(Vrms2/50/0.001) to solve for the voltage associated with 19 dBm.

Attachment 1: VCOOutputvsInput.png
VCOOutputvsInput.png
  197   Wed Jul 7 01:02:06 2010 taraDailyProgressLaserLaser driver is stationed on the electronic rack

I moved the laser driver to the electronic rack and it should stay there for the final setup.

A cable for interlock is made and connected to the laser driver.

A cable for slow actuator is also made and connected. Now we can use the medm FSS screen instead of the voltage calibrator

to adjust the NPRO's temperature (slow actuator.)

 

I went to 40m lab for a measurement

 

 

  196   Fri Jul 2 13:45:55 2010 FrankNotesEnvironmentchanged lab temperature

lowered the lab temperature avout 1K to get some data for the time constant from the environment to the chamber over the weekend which i can compare with my theoretical model.

current heater voltage for revcav is 14.5V @37Ohm resistance, so heater power is currently 5.68W.
current chamber temp is ~29C
current RT is 21.3C

  195   Thu Jul 1 13:49:58 2010 MeganNotesVCOVCO Output Amplitude Depending on Modlevel Input

I measured the output amplitude of the VCO with different Modlevel voltage inputs (from 0 V to 5 V) by measuring the RMS voltage on the oscilloscope. While I was varying the voltage input, I discovered the sine wave is very distorted between roughly 0.8 V and 3.5 V of modlevel input. Extra peaks started appearing that disappeared or became less prominent when below or above that range. The graphs show the output of the VCO in dBm and voltage (input impedance of 50 Ohms), accounting for the 19 dBm attenuation added to the output of the VCO.

I also started measuring the phase noise of the complete, unmodified VCO so we have a reference for the noise of the modified version when we are done with that. It is all set up right now and I should be able to start recording data on the computer soon. After we have a reference of the noise of the unmodified VCO, we will modify the VCO and see how it affects the noise level - hopefully it drops.

Attachment 1: Output_vs_Modlevel_Voltage.png
Output_vs_Modlevel_Voltage.png
  194   Wed Jun 30 23:26:27 2010 taracDailyProgressRefCavDebugging FSS servo+Locking RefCav

summary of this entry

  RefCav was not locked becase of the RFPD's input power. I haven't checked yet if it's the power supply or the cable. Now RefCav is locked.

 

Details

I was trying to lock RefCav again, but it didn't work.

So, I checked the error signal from mixer out, saw nothing.

It turned out that RF out from the PD has low voltage output, ~10mV.

I unplugged the cable and switch to another PD's cable along with its power supply, the DC increased to ~300 mV, a good sign.

Still, no error signal from Mixer Out Channel, nothing at all.

The control loop is not closed. Now 35.5 MHz LO is connected to the EOM, and to the servo card, LO input channel. The RF signal goes to the PD input.

We should be able to see the error signal from this setup.

**********The loop in the medm screen must be enabled in order to see the error signal******.

 

     EOM  -----------------BS--------------------- [ RefCav ]

        |                           |

        LO                       PD

        |                             |

        -------- X -------------

                   |

                   --->  Error signal

 

  I mix the signals from PD and LO by a Mixer [Minicurcuit ZFM 3 S+] and see something. It's a peak crossing zero at center instead of three peaks like the error signals. 

So I'm not sure what I'm seeing, but it seems that Mixer Out channel from the spare FSS servo might not working. I switched back to the first card, and now I see the error signals.

I set the RF to 6.1 V which corresponds to 97 mV of the error signal's P-P height.

Phase adj: is 4.5122 V,

phase flip: 0

Gain: [not set yet]

      All these values are saved in startup

  To summary, RefCav was not locked becase of the power cable which I haven't checked yet if it's the power supply or the cable. Now it's locked.

I'll set the gain tomorrow and see where I can connect the slow actuator for the laser driver. Right now I can't find where the servo for slow actuator would be.

 

ELOG V3.1.3-